A bicycle travels up a hill at an average speed of 4m/s and down the same hill through the same distance at an average speed of 8m/s. Express the average speed for the entire journey in terms oftandd.

Grade:12

1 Answers

raj
10 years ago
There is nothing difficult in this question. It gives you a lucid description of the situation and the values are given conspicuously. All you need to do is to understand thatF x s= work done, that is, represented by the letterEin the formula: P = E/t. The rest of the solution is given below: P = E/t = (F x s)/ t = F x (s/t) = F x speed = 15N x 0.5 m/min = 0.125W